# The hereditariness problem for the Černý conjecture

### v1

@@ 7.2 The semisimple case
We now address the hereditariness problem for semisimple automata. We begin with the study of the relationship between congruences and ideals for semisimple but non-simple automata. In particular, we relate the minimal ideals of $\mathcal{A}^{*}$ to the minimal non-trivial elements of $\mathrm{Cong}(\mathcal{A})$ via our Galois connection.
@@
Let $\mathcal{A}$ be a synchronizing DFA and $I_{1},I_{2}\subseteq\mathcal{A}^{*}$ be two distinct non-trivial minimal ideals. Then $\rho(I_{1}),\rho(I_{2})$ are non-trivial congruences of $\mathcal{A}$. In particular, if $\mathcal{A}$ is simple then $\mathcal{A}^{*}$ admits a unique minimal ideal.
@@
Let $\mathcal{A}$ be a semisimple DFA such that $\sigma\in\text{Cong}(\mathcal{A})\setminus\{\Delta_{\mathcal{A}},\nabla_{\mathcal{A}}\}$ is a non-trivial minimum, and $I(\sigma)\neq\text{Syn}(\mathcal{A})$. Then $\text{Syn}(\mathcal{A})=\text{Syn}(\mathcal{A}/\sigma)$. In particular, $\mathcal{A}/\sigma$ is semisimple.
@@
To reach a contradiction assume that $\text{Syn}(\mathcal{A}/\sigma)\neq\text{Syn}(\mathcal{A})$. Since $\text{Syn}(\mathcal{A}),I(\mathcal{A})\neq\text{Syn}(\mathcal{A})$, in $\mathcal{A}^{*}$ there are minimal non-trivial ideals $I_{1},I_{2}$ with $I_{1}\subseteq\text{Syn}(\mathcal{A}/\sigma),\ I_{2}\subseteq I(\sigma)$. With abuse of notation, we consider $I_{1},I_{2}$ as ideals of $\text{M}(\mathcal{A})$ by means of the Rees morphism $\theta:\text{M}(\mathcal{A})\rightarrow\mathcal{A}^{*}$, and using $I_{i}$ also to refer to $\theta^{-1}(I_{i})$. Assume that $I_{1}=I_{2}$: by Lemma 5.4 we would have that
$I_{1}\subseteq\text{Syn}(\mathcal{A}/\sigma)\cap I(\sigma)\subseteq\text{Rad}(\mathcal{A}^{*})=\text{Syn}(\mathcal{A})$
which is clearly a contradiction. Assume then $I_{1}\neq I_{2}$ and observe that from Lemma 7.7 together with the fact that $\sigma$ is minimum, we get $\rho(I_{1})\supseteq\sigma$. Thus, by means of the Galois Connection we have $I_{1}\subseteq I(\rho(I_{1}))\subseteq I(\sigma).$ But then $I_{1}\subseteq\text{Syn}(\mathcal{A}/\sigma)\cap I(\sigma)$, hence by Lemma 5.4 together with the semisimplicity of $\mathcal{A}$ we get $I_{1}\subseteq\text{Syn}(\mathcal{A})$, which is a contradiction. Hence, we may conclude that $\text{Syn}(\mathcal{A}/\sigma)=\text{Syn}(\mathcal{A})$ holds.
@@
A synchronizing DFA $\mathcal{A}$ is said to be quasi-simple if $\text{Cong}(\mathcal{A})$ admits a nontrivial minimum $\sigma$ and $I(\sigma)=\text{Syn}(\mathcal{A})$.
@@
Let $\sigma\in\text{Cong}(\mathcal{A})$ be the non-trivial minimum and $\rho:=\rho(\text{Rad}(\mathcal{A}))$. Assume $\mathcal{A}$ non-semisimple, by Lemma  6.1, $\rho$ is non-trivial and thus $\sigma\subseteq\rho$. From this and the property of the Galois connection, we obtain:
@@
Observe that the only non-trivial congruence is $\sigma:=\{\{q_{1}\},\{q_{2},q_{3}\}\}$ and $\mathcal{A}$ semisimple with $I(\sigma)=\text{M}(\mathcal{A})\pi(a)\text{M}(\mathcal{A})=\text{Syn}(\mathcal{A})$.
@@
Let $\mathcal{A}$ be a quasi-simple DFA and $\sigma\in\text{Cong}(\mathcal{A})$ be its minimum congruence. Is it true that $\mathcal{A}/\sigma$ is semisimple?
@@
Let $\mathcal{A}$ be a quasi-simple automaton, and let $\sigma\in\text{Cong}(\mathcal{A})$ denote its unique non-trivial minimal congruence. Then, for every non-trivial $\tau\in\text{Cong}(\mathcal{A})$, we have $I(\tau)=\text{Syn}(\mathcal{A})$. Indeed, by the Galois connection,
@@
which forces $I(\tau)=\text{Syn}(\mathcal{A})$. Furthermore, if $I\neq\text{Syn}(\mathcal{A})$ is an ideal of $M(\mathcal{A})$, then $\rho(I)=\Delta_{\mathcal{A}}$. Together, these facts show that the only fixed points of the Galois connection are $\text{Syn}(\mathcal{A})$ and $M(\mathcal{A})$ in $\mathcal{I}(\mathcal{A})$, and $\Delta_{\mathcal{A}}$ and $\nabla_{\mathcal{A}}$ in $\text{Cong}(\mathcal{A})$.
@@
Let $\mathcal{A}$ be a simple or quasi-simple DFA. Then $\mathcal{A}^{*}$ admits a unique $0$-minimal ideal.
@@
If $\mathcal{A}$ is simple, we are done by Lemma 7.7. Assume then $\mathcal{A}$ to be quasi-simple. Let $I_{1},I_{2}\subseteq\mathcal{A}^{*}$ be two minimal non-zero ideals and assume $I_{1}\neq I_{2}$. Again by Lemma 7.7, we have $\rho(I_{2})\neq\Delta_{\mathcal{A}}$: again if $\sigma$ is the minimum congruence we get $I_{2}\subseteq I(\sigma)=0$ and thus we can conclude that $I_{2}=0$, which is a contradiction. ∎
@@
Let $\mathcal{A}$ be a semisimple DFA which is not simple, and let $\mathcal{C}=\{\sigma_{1},\ldots,\sigma_{s}\}\subseteq\text{Cong}(\mathcal{A})$ be the set of its non-trivial minimal congruences. Assume that $s\geq 2$ and $\text{Syn}(\mathcal{A}/\sigma)\neq\text{Syn}(\mathcal{A})$ for every $\sigma\in\mathcal{C}$. Then, there exist distinct $0$-minimal ideals $I_{1},\ldots,I_{s}$ in $\mathcal{A}^{*}$ such that for every $i\in[1,s]$ we have:
@@
Hence, $(p,q)\in\rho(I_{j})\cap\sigma_{i}$, so in particular $\rho(I_{j})\cap\sigma_{i}\neq\Delta_{\mathcal{A}}$. By the minimality of $\sigma_{i}$, it follows that $\sigma_{i}\subseteq\rho(I_{j})$. Now, by Remark 5.6 and the inclusion $\sigma_{i}\subseteq\rho(I_{j})$, we deduce that
@@
  • If $\text{Cong}(\mathcal{A})$ admits a non-trivial minimum $\sigma$, then by Proposition 7.8 we have that $\mathcal{A}/\sigma$ is semisimple and $\text{Syn}(\mathcal{A})=\text{Syn}(\mathcal{A}/\sigma)$. Then by induction hypothesis we have that $\exists u\in\text{Syn}(\mathcal{A}/\sigma)=\text{Syn}(\mathcal{A})$ such that $|u|\leq(n-2)^{2}<(n-1)^{2}$.
  • Assume that $\mathcal{A}$ admits $k$-minimal non-trivial atoms, namely $\mathcal{C}=\{\sigma_{1},\ldots,\sigma_{k}\}$ with $k>1$. If $\text{Syn}(\mathcal{A}/\sigma_{i})=\text{Syn}(\mathcal{A})$ for some $i\in\{1,\ldots,k\}$, then by Lemma 4.9 the quotient $\mathcal{A}/\sigma_{i}$ is semisimple. By the induction hypothesis, there exists $u\in\text{Syn}(\mathcal{A}/\sigma_{i})=\text{Syn}(\mathcal{A})$ such that $|u|\leq(n-2)^{2}$, and we are done. Hence, we may assume that $\text{Syn}(\mathcal{A}/\sigma_{i})\neq\text{Syn}(\mathcal{A})$ for every $i$. Consider the congruences
@@
    where $\tau_{i}=\rho(I_{i})$ as defined in Lemma 7.16. By that lemma, we have $\tau_{i}\supseteq\sigma_{k}$ for every $i\in\{1,\ldots,k-1\}$, and thus $\tau\neq\Delta_{\mathcal{A}}$. Furthermore, Proposition 4.11 combined with Lemma 7.18 implies that $\mathcal{A}/\tau$ is semisimple; the same lemma also guarantees that $\mathcal{A}/\rho=\mathcal{A}/\tau_{k}$ is semisimple. By the induction hypothesis, let $u\in\text{Syn}(\mathcal{A}/\rho)$ and $v\in\text{Syn}(\mathcal{A}/\tau)$ be reset words satisfying the Černý conjecture in their respective quotients. We have $Q\cdot v\subseteq[p]_{\tau}$ for some $p\in Q$. Assume that $\tau$ admits a $1$-class, i.e., $q]_{\tau}=\{q\}$ for some $q\in Q$. By the strong connectedness of $\mathcal{A}/\tau$ (which comes from the same property for $\mathcal{A}$), there exists $z\in\Sigma^{*}$ with $|z|\leq n-2$ such that $p]_{\tau}\cdot z=[q]_{\tau}=\{q\}$. Hence $vz\in\text{Syn}(\mathcal{A})$ with
@@
We conclude this work by presenting a diagram depicted in Fig. 1 showing how the different classes of automata introduced in this paper are related by inclusion. We also included references to examples that show these inclusions are proper and non-trivial. The only missing example (which, as far as we know, is still unknown) is that of a simple but non-irreducible automaton, which suggests that these two classes might actually coincide.

### v2

@@ 7.2 The semisimple case
We now address the hereditariness problem for semisimple automata. We begin with the study of the relationship between congruences and ideals for semisimple but non-simple automata. In particular, we relate the minimal ideals of $\mathcal{A}^{\star}$ to the minimal non-trivial elements of $\mathrm{Cong}(\mathcal{A})$ via our Galois connection.
@@
Let $\mathcal{A}$ be a synchronizing DFA and $I_{1},I_{2}\subseteq\mathcal{A}^{\star}$ be two distinct 0-minimal ideals. Then $\rho(I_{1}),\rho(I_{2})$ are non-trivial congruences of $\mathcal{A}$. In particular, if $\mathcal{A}$ is simple then $\mathcal{A}^{\star}$ admits a unique 0-minimal ideal.
@@
Let $\mathcal{A}$ be a semisimple DFA admitting a monolith $\sigma$, and $I(\sigma)\neq\text{Syn}(\mathcal{A})$. Then $\text{Syn}(\mathcal{A})=\text{Syn}(\mathcal{A}/\sigma)$. In particular, $\mathcal{A}/\sigma$ is semisimple.
@@
To reach a contradiction assume that $\text{Syn}(\mathcal{A}/\sigma)\neq\text{Syn}(\mathcal{A})$. Since $\text{Syn}(\mathcal{A}/\sigma),I(\sigma)\neq\text{Syn}(\mathcal{A})$, in $\mathcal{A}^{\star}$ there are minimal non-trivial ideals $I_{1},I_{2}$ with $I_{1}\subseteq\text{Syn}(\mathcal{A}/\sigma),\ I_{2}\subseteq I(\sigma)$. With abuse of notation, we consider $I_{1},I_{2}$ as ideals of $\text{M}(\mathcal{A})$ by means of the Rees morphism $\theta:\text{M}(\mathcal{A})\rightarrow\mathcal{A}^{\star}$, and using $I_{i}$ also to refer to $\theta^{-1}(I_{i})$. Assume that $I_{1}=I_{2}$: by Lemma 5.4 we would have that
$I_{1}\subseteq\text{Syn}(\mathcal{A}/\sigma)\cap I(\sigma)\subseteq\text{Rad}(\mathcal{A}^{\star})=\text{Syn}(\mathcal{A})$
which is clearly a contradiction. Assume then $I_{1}\neq I_{2}$ and observe that from Lemma 7.7 together with the fact that $\sigma$ is monolith, we get $\rho(I_{1})\supseteq\sigma$. Thus, by means of the Galois Connection we have $I_{1}\subseteq I(\rho(I_{1}))\subseteq I(\sigma).$ But then $I_{1}\subseteq\text{Syn}(\mathcal{A}/\sigma)\cap I(\sigma)$, hence by Lemma 5.4 together with the semisimplicity of $\mathcal{A}$ we get $I_{1}\subseteq\text{Syn}(\mathcal{A})$, which is a contradiction. Hence, we may conclude that $\text{Syn}(\mathcal{A}/\sigma)=\text{Syn}(\mathcal{A})$ holds.
@@
A synchronizing DFA $\mathcal{A}$ is said to be quasi-simple if $\mathcal{A}$ admits a monolith $\sigma$ and $I(\sigma)=\text{Syn}(\mathcal{A})$.
@@
Let $\sigma$ be the monolith of $\mathcal{A}$ and $\rho:=\rho(\text{Rad}(\mathcal{A}))$. Assume $\mathcal{A}$ non-semisimple, by Lemma  6.1, $\rho$ is non-trivial and thus $\sigma\subseteq\rho$. From this and the property of the Galois connection, we obtain:
@@
Observe that the only non-trivial congruence is $\sigma:=\{\{q_{1}\},\{q_{2},q_{3}\}\}$ and $\mathcal{A}$ is semisimple with $I(\sigma)=\text{M}(\mathcal{A})\pi(a)\text{M}(\mathcal{A})=\text{Syn}(\mathcal{A})$.
@@
Let $\mathcal{A}$ be a quasi-simple DFA and $\sigma\in\text{Cong}(\mathcal{A})$ be its monolith. Is it true that $\mathcal{A}/\sigma$ is semisimple?
@@
Let $\mathcal{A}$ be a quasi-simple automaton, and let $\sigma\in\text{Cong}(\mathcal{A})$ denote its unique non-trivial minimal congruence. Then, for every non-trivial $\tau\in\text{Cong}(\mathcal{A})$, we have $I(\tau)=\text{Syn}(\mathcal{A})$. Indeed, by the Galois connection we have:
@@
which forces $I(\tau)=\text{Syn}(\mathcal{A})$. Furthermore, if $I\neq\text{Syn}(\mathcal{A})$ is an ideal of $\text{M}(\mathcal{A})$, then $\rho(I)=\Delta_{\mathcal{A}}$. Together, these facts show that the only fixed points of the Galois connection are $\text{Syn}(\mathcal{A})$ and $\text{M}(\mathcal{A})$ in $\mathcal{I}(\mathcal{A})$, and $\Delta_{\mathcal{A}}$ and $\nabla_{\mathcal{A}}$ in $\text{Cong}(\mathcal{A})$.
@@
Let $\mathcal{A}$ be a simple or quasi-simple DFA. Then $\mathcal{A}^{\star}$ admits a unique 0-minimal ideal.
@@
If $\mathcal{A}$ is simple, we are done by Lemma 7.7. Assume then $\mathcal{A}$ to be quasi-simple. Let $I_{1},I_{2}\subseteq\mathcal{A}^{\star}$ be two minimal non-zero ideals and assume $I_{1}\neq I_{2}$. Again by Lemma 7.7, we have $\rho(I_{2})\neq\Delta_{\mathcal{A}}$: again if $\sigma$ is the monolith we get $I_{2}\subseteq I(\sigma)=0$ and thus we can conclude that $I_{2}=0$, which is a contradiction. ∎
@@
Let $\mathcal{A}$ be a semisimple DFA which is not simple, and let $\mathcal{C}=\{\sigma_{1},\ldots,\sigma_{s}\}\subseteq\text{Cong}(\mathcal{A})$ be the set of its non-trivial minimal congruences. Assume that $s\geq 2$ and $\text{Syn}(\mathcal{A}/\sigma)\neq\text{Syn}(\mathcal{A})$ for every $\sigma\in\mathcal{C}$. Then, there exist distinct 0-minimal ideals $I_{1},\ldots,I_{s}$ in $\mathcal{A}^{\star}$ such that for every $i\in[1,s]$ we have:
@@
Hence, $(p,q)\in\rho(I_{j})\cap\sigma_{i}$, so in particular $\rho(I_{j})\cap\sigma_{i}\neq\Delta_{\mathcal{A}}$. By the minimality of $\sigma_{i}$, it follows that $\sigma_{i}\subseteq\rho(I_{j})$. Now, by Proposition 5.6 and the inclusion $\sigma_{i}\subseteq\rho(I_{j})$, we deduce that
@@
  • If $\mathcal{A}$ admits a monolith $\sigma$, then by Proposition 7.8 we have that $\mathcal{A}/\sigma$ is semisimple and $\text{Syn}(\mathcal{A})=\text{Syn}(\mathcal{A}/\sigma)$. Then by induction hypothesis we have that $\exists u\in\text{Syn}(\mathcal{A}/\sigma)=\text{Syn}(\mathcal{A})$ such that $|u|\leq(n-2)^{2}<(n-1)^{2}$.
  • Assume that $\mathcal{A}$ admits $k$ minimal non-trivial atoms, namely $\mathcal{C}=\{\sigma_{1},\ldots,\sigma_{k}\}$ with $k>1$. If $\text{Syn}(\mathcal{A}/\sigma_{i})=\text{Syn}(\mathcal{A})$ for some $i\in\{1,\ldots,k\}$, then by Lemma 4.5 together with the semisimplicity of $\mathcal{A}$ we have:
    $\text{Rad}(A/\sigma_{i})=\text{Rad}(\mathcal{A})=\text{Syn}(\mathcal{A})=\text{Syn}(\mathcal{A}/\sigma_{i})$
    and thus the quotient $\mathcal{A}/\sigma_{i}$ is semisimple. By the induction hypothesis, there exists $u\in\text{Syn}(\mathcal{A}/\sigma_{i})=\text{Syn}(\mathcal{A})$ such that $|u|\leq(n-2)^{2}$, and we are done. Hence, we may assume that $\text{Syn}(\mathcal{A}/\sigma_{i})\neq\text{Syn}(\mathcal{A})$ for every $i$. Consider the congruences
@@
    where $\tau_{i}=\rho(I_{i})$ as defined in Lemma 7.16. By that lemma, we have $\tau_{i}\supseteq\sigma_{k}$ for every $i\in\{1,\ldots,k-1\}$, and thus $\tau\neq\Delta_{\mathcal{A}}$. Furthermore, Proposition 4.7 combined with Lemma 7.18 implies that $\mathcal{A}/\tau$ is semisimple; the same lemma also guarantees that $\mathcal{A}/\rho=\mathcal{A}/\tau_{k}$ is semisimple. By the induction hypothesis, let $u\in\text{Syn}(\mathcal{A}/\rho)$ and $v\in\text{Syn}(\mathcal{A}/\tau)$ be a Černý reset words in their respective quotients. We have $Q\cdot v\subseteq[p]_{\tau}$ for some $p\in Q$. Assume that $\tau$ admits a $1$-class, i.e., $q]_{\tau}=\{q\}$ for some $q\in Q$. By the strong connectedness of $\mathcal{A}/\tau$ (which comes from the same property for $\mathcal{A}$), there exists $z\in\Sigma^{*}$ with $|z|\leq n-2$ such that $p]_{\tau}\cdot z=[q]_{\tau}=\{q\}$. Hence $vz\in\text{Syn}(\mathcal{A})$ with
@@
We conclude this work by presenting a diagram depicted in Fig. 1 showing how the different classes of automata introduced in this paper are related by inclusion. We also included references to examples that show these inclusions are proper and non-trivial. The only missing example (which, as far as we know, is still unknown) is that of a simple but non-irreducible automaton, which suggests that these two classes might actually coincide. We also believe that the class of simple (respectively, irreducible) automata has measure one with respect to the uniform distribution. More precisely, we pose the following conjecture.
Open Problem 7.22.
Let $n\in\mathbb{N}$, and let $\mathcal{A}_{n}$ be a deterministic finite automaton with $n$ states, chosen uniformly at random among all complete deterministic automata over a fixed finite alphabet of size greater than $1$. Is it true that the probability that $\mathcal{A}_{n}$ is simple tends to $1$ as $n\to\infty$? What can be said about the corresponding statement for irreducibility?
@@
Acknowledgments
We are grateful to the anonymous referees for their careful reading and valuable feedback, which helped us improve the clarity of the paper.